### v2

Abstract
We investigate the asymptotic symmetry structure of two–dimensional dilaton gravity in its $\mathcal{N}=1$ supersymmetric extension based on the $\mathfrak{osp}(1|2)$ Lie superalgebra. Within the BF theoretical framework, we analyze affine and superconformal boundary conditions for each case and systematically derive the associated asymptotic symmetry algebras. While the classical theory recovers the Virasoro algebra or its affine enhancement, the supersymmetric extension yields a classical $\mathcal{N}=1$ superconformal algebra, subject to dynamical symmetry breaking mechanisms induced by the dilaton supermultiplet. We find that the boundary behavior of the dilaton not only leads to a reduction of the full affine $\mathfrak{osp}(1|2)_{k}$ symmetry down to $\tt{O}\tt{S}p(1|2)$, but also introduces an abelian extension through commuting modes. These results reveal a novel interplay between symmetry breaking and symmetry extension in low-dimensional supergravity. Our construction generalizes previous analyses of $sl(2,\mathbb{R})$ dilaton gravity to the supersymmetric domain and offers a consistent foundation for investigating boundary dynamics beyond the Schwarzian regime.
Keywords:
Jackiw–Teitelboim(JT) supergravity, asymptotic symmetries, superconformal field theory, AdS/CFT, holographic duality
1 Introduction
The holographic principle Maldacena:1997re asserts that a gravitational theory defined in $d$ dimensions can be reformulated through an equivalent boundary theory living in (d1)(d{-}1) dimensions. This paradigm inspired investigations into two–dimensional $\mathrm{AdS}_{2}$ holography in the late 1990s, primarily driven by the analysis of asymptotic symmetries in Jackiw–Teitelboim (JT) gravity Teitelboim:1983ux ; Jackiw:1984je and dilaton models arising from dimensional reduction Strominger:1998yg ; Cadoni:1998sg ; Hotta:1998iq ; Navarro-Salas:1999zer . These studies have recently experienced a resurgence, motivated by the conjectured duality between JT gravity and the Sachdev–Ye–Kitaev (SYK) model Maldacena:2016hyu ; Jensen:2016pah ; Sachdev:1992fk ; Kitaev , which exemplifies a solvable system with quantum chaotic behavior. Within this correspondence, the Schwarzian boundary action emerges as a low-energy effective description Mertens:2018fds .
Conformal field theory has emerged as a foundational element in contemporary theoretical physics, with far–reaching applications from string theory to the study of complex many–body systems. Its significance lies in the stringent symmetry principles it enforces, which serve to tightly constrain the landscape of admissible physical models. In two dimensions, conformal invariance proves exceptionally potent, manifesting in the form of the infinite–dimensional Virasoro algebra Virasoro:1969zu . Over time, extensive research has been directed toward extending this symmetry to more generalized algebraic settings, such as the Kac–Moody algebras Kac_1968 ; Moody:1968zz and WW–algebras Zamolodchikov:1985wn . Simultaneously, substantial effort has been invested in deepening the theoretical understanding of the Virasoro structure itself DiFrancesco:1997nk ; Blumenhagen:2009zz ; Prochazka:2014gqa ; Ozer:2015dha . These algebraic symmetries also underpin the formulation of two–dimensional dilaton gravity, wherein the appearance of the Virasoro algebra at the boundary offers a natural bridge between conformal techniques and gravitational dynamics.
Two–dimensional pure gravity is identified as a topological theory, owing to the lack of propagating local degrees of freedom. In $\mathrm{AdS}_{2}$ geometries, the symmetry group is given by ${\tt S}{\tt L}(2,\mathbb{R})$, which functions both as the isometry and asymptotic symmetry group and implies a vanishing central charge Almheiri:2014cka . The incorporation of a dilaton field, however, renders the theory dynamical: the dilaton couples to the metric and activates additional boundary excitations, leading to frameworks such as Jackiw–Teitelboim gravity Teitelboim:1983ux ; Jackiw:1984je ; Maldacena:2016upp ; Jensen:2016pah . This modification extends the symmetry algebra from ${\tt S}{\tt L}(2,\mathbb{R})$ to the infinite–dimensional Virasoro algebra Cadoni:1999ja , thereby allowing for a nonvanishing central charge. JT gravity, most notably through its Schwarzian boundary term, has become a powerful tool for exploring aspects of quantum gravity, particularly in light of its holographic duality with the SYK model Maldacena:2016upp ; Engelsoy:2016xyb ; Almheiri:2014cka . These developments have placed two–dimensional dilaton gravity at the core of contemporary investigations into holography, black hole thermodynamics, and quantum information Cvetic:2016eiv ; Sachdev:2019bjn ; Sachdev:2010um .
The intricate internal structure of JT gravity is fundamentally grounded in its boundary symmetry content and the formal mathematical scaffolding that supports it. The theory maintains a profound connection to conformal symmetry through its interplay with the Virasoro algebra and Schwarzian derivatives Iliesiu ; Mertens ; OvisenkoSchwarzian , and it has been systematically broadened by the implementation of relaxed boundary conditions and higher–spin extensions Gross:2016kjj ; Witten:2016iux ; Yoon . The boundary behavior can also be effectively captured using coadjoint orbit methods, which serve as powerful tools for probing symmetry–breaking phenomena and their physical ramifications Davison:2016ngz ; NarayanYoon ; Fu:2016vas ; Grumiller:2017qao ; Alkalaev:2013fsa ; Grumiller:2013swa ; Grumiller ; Kine ; Fukuyama . While the Schwarzian formalism remains a crucial link to the SYK dual, the scope of asymptotic symmetry analysis extends beyond it; in particular, the BF formulation of JT gravity offers a structurally independent yet equally robust avenue for such inquiries Grumiller:2002nm ; Banks:1990mk . Within this formalism, the dilaton field functions as a Lagrange multiplier that enforces curvature constraints, and the associated asymptotic symmetry algebra can be extended, for instance, to a Virasoro–Kac–Moody structure, via the inclusion of internal symmetry sectors Ikeda:1993fh ; Jackiw:1982hg . This flexibility underscores the utility of JT gravity as a versatile platform for studying gravitational dynamics in lower dimensions, irrespective of adherence to the Schwarzian perspective. This methodological structure aligns with the strategy developed in our previous work on bosonic $\mathfrak{sl}(3,\mathbb{R})$ JT gravity Ozer:2025bpb , and here we implement its supersymmetric analogue using the $\mathfrak{osp}(1|2)$ superalgebra. While prior analyses of JT supergravity, particularly those centered around the Schwarzian action Almheiri:2014cka ; Maldacena:2016upp ; Cadoni:1999ja ; Engelsoy:2016xyb ; Cvetic:2016eiv ; Sachdev:2019bjn ; Sachdev:2010um , focus on the boundary effective theory and its relation to the SYK model, our approach emphasizes the bulk gauge-theoretic formulation and its implications for asymptotic symmetry breaking. In contrast to the boundary-centered derivations of the ${\cal N}=1$ super–Schwarzian theory, we construct the full $\mathfrak{osp}(1|2)$ asymptotic algebra via BF theory and coadjoint orbit methods, thereby uncovering dynamical constraints absent in Schwarzian-only treatments. This strategy is conceptually aligned with our earlier $\mathfrak{sl}(3,\mathbb{R})$ construction Ozer:2025bpb , and extends the findings of Grumiller:2002nm ; Banks:1990mk ; Ikeda:1993fh ; Jackiw:1982hg into the supersymmetric domain. As such, the present work provides a structurally distinct extension of JT supergravity beyond existing formulations.
In the framework of two–dimensional JT gravity, the Schwarzian formalism is frequently utilized to characterize boundary dynamics and to establish its correspondence with the SYK model Grumiller:2002nm ; Ikeda:1993fh ; Jackiw:1982hg . Nevertheless, when the objective shifts toward understanding the structure of asymptotic symmetries rather than boundary–specific features, reliance on the Schwarzian approach becomes nonessential Grumiller:2002nm ; Jackiw:1982hg . In analogy with the Chern–Simons treatment of three–dimensional gravity and the associated Brown–Henneaux boundary conditions, JT gravity admits a BF–theoretic formulation where the dilaton functions as a Lagrange multiplier enforcing curvature constraints Grumiller:2002nm ; Banks:1990mk . This perspective enables the realization of extended asymptotic symmetry algebras, such as the Virasoro or Virasoro–Kac–Moody structures, via the inclusion of internal symmetries and the imposition of compatible boundary data Ikeda:1993fh ; Jackiw:1982hg . In particular, coadjoint orbit techniques Ikeda:1993fh ; Jackiw:1982hg provide a systematic route for exploring such algebras, especially in scenarios featuring boundary–induced symmetry breaking. As a result, the BF formulation accentuates the adaptability of JT gravity in addressing holographic and asymptotic structures beyond the confines of the Schwarzian perspective.
The structural sophistication of JT gravity stems from the intricate interaction between its boundary symmetry content and formal mathematical architecture. The model exhibits a deep connection to conformal symmetry through the Virasoro algebra and Schwarzian derivatives Iliesiu ; Mertens ; OvisenkoSchwarzian , and its scope has been broadened by incorporating relaxed boundary conditions and higher–spin enhancements Gross:2016kjj ; Witten:2016iux ; Yoon . Within this setting, coadjoint orbit techniques provide a powerful framework for probing symmetry breaking phenomena and their physical ramifications Davison:2016ngz ; NarayanYoon ; Fu:2016vas ; Grumiller:2017qao ; Alkalaev:2013fsa ; Grumiller:2013swa ; Grumiller ; Kine ; Fukuyama . Although the Schwarzian approach plays a central role in the SYK duality, the study of asymptotic symmetries does not intrinsically depend on it Grumiller:2002nm ; Banks:1990mk . The BF formulation is particularly well-suited in this regard, as it seamlessly integrates both affine and conformal symmetry realizations Ikeda:1993fh ; Jackiw:1982hg . In the present study, we expand this line of investigation into the realm of $\mathfrak{osp}(1|2)$ supergravity, aiming to construct supersymmetric asymptotic structures that extend beyond the Schwarzian regime.
[…]
Within AdS2 gravity, boundary conditions are generally classified into two dominant types: the broader affine class and the more constrained conformal class. Each class is associated with a distinct asymptotic symmetry algebra and yields unique boundary dynamics. The affine category naturally arises in the Poisson sigma model, where both the dilaton and the boundary metric are allowed to fluctuate at leading order. This framework gives rise to an infinite–dimensional centerless affine $\mathfrak{sl}(2,\mathbb{R})$ current algebra. While the Schwarzian action can be seen as a projection embedded within this more general construction, the complete theory supports a considerably richer algebraic structure. On the other hand, conformal boundary conditions impose geometric constraints that lead to a low–energy effective Schwarzian theory, describing pseudo–Goldstone modes resulting from spontaneous breaking of reparametrization symmetry. These configurations exhibit an off–shell Virasoro symmetry that reduces to ${\tt S}{\tt L}(2,\mathbb{R})$ on-shell, and are widely regarded as the gravitational duals to the infrared limit of the SYK model. Altogether, the affine and conformal classes of boundary conditions define complementary sectors in AdS2 holography and highlight the conceptual breadth of two–dimensional dilaton gravity beyond the confines of the JT framework.
Unlike the well–established $sl(2,\mathbb{R})$ formulation of JT gravity, where boundary symmetries are governed by the linear Virasoro algebra, the $sl(3,\mathbb{R})$ extension offers a fertile ground for exploring more complex, nonlinear structures such as the $\mathcal{W}_{3}$ algebra Ozer:2025bpb . This framework is motivated by the pursuit of higher-spin generalizations of dilaton gravity and by the goal of understanding their impact on boundary dynamics. In particular, the inclusion of spin–3 fields alters the asymptotic symmetry structure: it not only enhances the existing symmetry but also introduces a tunable mechanism of symmetry breaking. This formulation yields fresh perspectives on how two–dimensional gravity responds to generalized boundary conditions and contributes to the broader effort to connect higher–spin JT gravity to SYK–like holographic duals. Furthermore, given that $\mathcal{W}_{3}$ algebras naturally arise in large–$N$ matrix models and in certain deformed SYK constructions, our setup may provide a plausible dual description for such deformed CFT1 theories.
Recent developments have employed JT gravity models augmented with ${\tt S}{\tt L}(3,\mathbb{R})$ symmetry to explore the influence of nonlinear structures, such as the $\mathcal{W}_{3}$ algebra, on boundary dynamics Ozer:2025bpb . These investigations have clarified how higher-spin fields deform the underlying asymptotic symmetry algebra and support symmetry-breaking mechanisms. In this work, we extend the analysis beyond the purely bosonic realm by introducing a supersymmetric generalization grounded in $\mathfrak{osp}(1|2)$ supergravity. In this framework, the incorporation of a dilaton supermultiplet leads to a novel family of asymptotic field configurations. This setup not only reconfigures the algebraic structure but also furnishes a natural platform for studying supersymmetry breaking patterns and their implications for boundary dynamics.
While the $\mathfrak{sl}(2,\mathbb{R})$ JT gravity model induces a truncation of the infinite–dimensional Virasoro algebra to its ${\tt S}{\tt L}(2,\mathbb{R})$ subgroup due to the boundary dynamics of the dilaton, the extended framework presented here carries out an analogous symmetry reduction for the $\mathcal{N}=1$ superconformal algebra. In this supersymmetric context, the dilaton and higher-spin fields jointly serve as dynamical stabilizers that restrict admissible asymptotic gauge variations, thereby enforcing a reduction of the full $\mathcal{N}=1$ structure to its $\mathfrak{osp}(1|2)$ subalgebra. This symmetry breaking not only produces a consistent algebraic truncation but also provides a coherent method for analyzing such reductions within higher–spin holography through the BF theory formalism. Throughout this analysis, the $\mathfrak{osp}(1|2)$ subalgebra is understood as the extended realization induced from the affine $\mathfrak{osp}(1|2)_{k}$ or $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ superconformal algebra, shaped by the dynamical influence of the dilaton sector.
This work centers on the extension of the asymptotic symmetry algebra arising from symmetry–breaking processes, with a particular focus on the $\mathfrak{osp}(1|2)$ generalization of the JT gravity framework. This setting enables a detailed exploration of how the algebraic structure evolves and how these modifications influence the underlying physical dynamics. Previous studies have demonstrated that while a Virasoro algebra can be reconstructed within two–dimensional dilaton gravity Grumiller:2013swa , the associated conserved charges were not entirely integrable, posing challenges for the formulation of a consistent AdS2/CFT1 correspondence. This issue was subsequently resolved by introducing dilaton–dependent charges and relaxing the assumption of Casimir invariance Grumiller:2017qao , which allowed for the full manifestation of the Virasoro symmetry and reinforced the conformal field theory interpretation. These advancements highlight the critical function of extended boundary conditions in AdS2 gravity and pave the way for deeper inquiries into its holographic structure.
In conclusion, this study undertakes a comprehensive investigation into the relationship between two–dimensional dilaton gravity, boundary conditions, and symmetry structures, aiming to fill key gaps identified in the existing body of research. The structural richness and symmetry features of JT gravity offer new angles from which to approach our understanding of holographic dualities. Within this broader landscape, two–dimensional gravity models have established themselves as foundational elements in theoretical physics, and the insights developed here are anticipated to inform and inspire subsequent explorations.
This paper investigates the asymptotic symmetry structure emerging in a two–dimensional JT supergravity model formulated on the foundation of the Lie superalgebra $\mathfrak{osp}(1|2)$. In this construction, the dilaton field is promoted from a classical scalar to a supersymmetry–sensitive supermultiplet, allowing not only reparametrization symmetries but also their supersymmetric counterparts to act nontrivially at the boundary. Our methodology adopts a BF theory formulation, wherein boundary conditions are leveraged to construct and systematically regulate the breaking of these extended supersymmetries. Anchored in coadjoint orbit techniques, this framework provides a firm basis for realizing supersymmetric holographic dualities that transcend the traditional Schwarzian setting.
The interplay between two–dimensional dilaton gravity and SYK–like systems has recently attracted renewed scholarly focus. In particular, the study presented in Momeni:2024ygv introduces a deformed SYK model as a plausible holographic dual to generalized JT gravity, contending that altered boundary terms and higher-spin extensions may serve to encode new structural elements of the corresponding dual theory.
In a complementary vein, multiple studies, such as Momeni:2020tyt ; Momeni:2020zkx ; Momeni:2021jhk , have investigated higher–spin extensions of AdS2 gravity, highlighting the central significance of extended symmetry algebras, especially $\mathcal{W}$–algebras, in organizing asymptotic dynamics. The present framework resonates with these contributions by formulating an $\mathfrak{osp}(1|2)$ extension of JT gravity within a BF–theoretic scheme that accommodates both symmetry breaking mechanisms and coadjoint orbit methodology. As a result, our construction yields a tangible realization of generalized boundary dynamics that naturally integrates into the broader context of higher–spin gravitational theories.
In this regard, the present construction is best interpreted as a supersymmetric extension of our previous bulk–based asymptotic symmetry analysis, rather than as a novel dynamical theory beyond the Schwarzian paradigm.
Our construction differs conceptually and technically from conventional JT gravity formulations based on $\mathfrak{sl}(2,\mathbb{R})$ Momeni:2024ygv ; Momeni:2020tyt ; Momeni:2020zkx , particularly those that emphasize boundary actions, dilaton constraints, or effective Schwarzian descriptions. While those models often encode symmetry breaking through boundary terms and thermodynamic consistency conditions, our approach derives asymptotic symmetries directly from the bulk via the structure of the gauge field and its residual supersymmetries. This method bypasses the need for an explicit boundary action, allowing a direct algebraic characterization of the asymptotic symmetry algebra. In that sense, our work aligns more closely with generalized bulk–based symmetry analyses, such as our earlier $\mathfrak{sl}(3,\mathbb{R})$ study Ozer:2025bpb , and represents a supersymmetric extension of that framework. We do not aim to construct a complete dual quantum mechanical theory; rather, we focus on the intrinsic gauge–theoretic origin of boundary symmetries in the supersymmetric BF setting.
Unlike earlier JT models based on $\mathfrak{sl}(2,\mathbb{R})$, where symmetry breaking typically appears through boundary dilaton constraints or horizon regularity conditions, our formulation realizes symmetry reduction algebraically via residual bulk supersymmetry, without imposing explicit dynamical constraints or thermodynamic inputs. On the physical level, symmetry breaking happens dynamically due to the time-dependent dilaton supermultiplet in the bulk, not because of boundary thermodynamics.
The structure of this paper is as follows: Section 2 introduces the quantum $\mathcal{N}=1$ superconformal algebra. It begins with an overview of the conformal structure and continues with the construction of its extended version incorporating conformal spins $(\Delta,1{-}\Delta)$. Section 3 develops the holographic dictionary for $sl(2,\mathbb{R})$ JT dilaton gravity. It starts from the BF theory formulation and elaborates on its boundary realizations, including both affine and conformal boundary conditions. Section 4 focuses on the $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ extension of higher–spin dilaton gravity. It analyzes the asymptotic symmetry algebra under affine boundary conditions and then presents the superconformal boundary structure that emerges in the Drinfeld–Sokolov gauge. Section 5 summarizes the main results and discusses their theoretical and holographic significance. Section 6 acknowledges the contributors and funding sources. Finally, Appendix A provides a summary of notation, while Appendix B reviews the boundary theories and their associated symmetry structures.
2 The quantum ${\cal N}=1$ superconformal algebra
We begin by summarizing the key structural elements of the quantum ${\cal N}=1$ superconformal algebra, focusing on the aspects most relevant to the framework developed in this work. Rather than presenting a complete treatment, our goal is to emphasize the defining features that inform the supersymmetric extensions we explore. Originally formulated by A.B. Zamolodchikov Zamolodchikov:1985wn , the quantum ${\cal N}=1$ superconformal algebra introduces a non–linear generalization of the Virasoro algebra that naturally accommodates higher–spin currents.
2.1 Conformal structure of the quantum ${\cal N}=1$ superconformal algebra
The algebra is built upon two primary fields: a fermionic current $G(z)$ of spin-$\tfrac{3}{2}$ and a bosonic stress–energy tensor $T(z)$ of spin-$2$. These fields admit standard mode expansions, given by $T(z)=\sum_{n}L_{n}z^{-n-2}$ and $G(z)=\sum_{n}G_{n}z^{-n-3/2}$. The essential structure of the quantum ${\cal N}=1$ superconformal algebra is encoded in their non–trivial operator product expansions, which reflect the characteristic non–linearity of this supersymmetric extension of the Virasoro algebra.
$T(z_{1})T(z_{2})$ $\sim\frac{\frac{c}{2}}{z_{12}^{4}}+\frac{2T}{z_{12}^{2}}+\frac{\partial T}{z_{12}},$ (1)
$T(z_{1})G(z_{2})$ $\sim\frac{\frac{3}{2}G}{z_{12}^{2}}+\frac{\partial G}{z_{12}},$ (2)
$G(z_{1})G(z_{2})$ $\sim\frac{\frac{2c}{3}}{z_{12}^{3}}+\frac{2T}{z_{12}}.$ (3)
[…]
In this section, we construct an extended version of the quantum ${\cal N}=1$ superconformal algebra by supplementing the conformal fields of spin $\Delta=2$ and $\Delta=3/2$, introduced in the previous subsection, with two additional components. These new fields, labeled $X$ and $Y$, are interpreted as dilaton partners and carry conformal weights $1-\Delta=-1$ and 1/2-1/2, respectively. Their standard mode expansions are given by $X(z)=\sum_{n}X_{n}z^{-n+1}$ and $F(z)=\sum_{n}F_{n}z^{-n+1/2}$. Using Thielemans’ Mathematica package Thielemans:1991uw , one obtains the non–vanishing operator product expansions that define the extended ${\cal N}=1$ superconformal algebra in this formulation.
$T(z_{1})T(z_{2})$ $\sim\frac{\frac{c}{2}}{z_{12}^{4}}+\frac{2T}{z_{12}^{2}}+\frac{\partial T}{z_{12}},$ (4)
$T(z_{1})G(z_{2})$ $\sim\frac{\frac{3}{2}G}{z_{12}^{2}}+\frac{\partial G}{z_{12}},$ (5)
$G(z_{1})G(z_{2})$ $\sim\frac{\frac{2c}{3}}{z_{12}^{3}}+\frac{2T}{z_{12}},$ (6)
$T(z_{1})X(z_{2})$ $\sim-\frac{X}{z_{12}^{2}}+\frac{\partial X}{z_{12}},$ (7)
$T(z_{1})F(z_{2})$ $\sim-\frac{\frac{1}{2}F}{z_{12}^{2}}+\frac{\partial F}{z_{12}},$ (8)
$G(z_{1})X(z_{2})$ $\sim-\frac{2F}{{z_{12}}},$ (9)
$G(z_{1})F(z_{2})$ $\sim-\frac{X}{{z_{12}^{2}}}+\frac{\partial X}{2z_{12}}.$ (10)
This extended version of the quantum ${\cal N}=1$ superconformal algebra will later be contrasted with its classical counterpart, which will be explicitly derived in Section 4.
3 The $sl(2,\mathbb{R})$ holographic dictionary
In this section, we explore the formalism of two–dimensional dilaton gravity defined on an $AdS_{2}$ background. We begin by outlining the essential features of dilaton gravity that are instrumental for constructing the holographic dictionary. This is followed by a systematic derivation of the asymptotic symmetry algebra within the $sl(2,\mathbb{R})$ framework. We then turn our attention to the role of external source terms and the resulting holographic Ward identities, which provide key insights into the structure of higher–spin black hole configurations.
3.1 JT dilaton gravity as a BF theory in two dimensions
This section provides a concise overview of higher–spin gravity in $AdS_{2}$, formulated within the framework of dilaton gravity. We adopt a gauge-theoretic perspective, treating dilaton gravity as a non–abelian theory. Specifically, our analysis focuses on the formulation of $AdS_{2}$ gravity in terms of the underlying $sl(2,\mathbb{R})$ algebraic structure.
3.2 Connection to $sl(2,\mathbb{R})$ BF theory
[…]
$S[{\cal X},{\cal A}]=\frac{k}{4\pi}\int_{{\cal M}}\mathfrak{tr}[{\cal X}{\cal F}]+S_{\rm bdy},$ (11)
[…]
The gauge fields $\cal A$} and $\cal X$} both take values in the Lie algebra $sl(2,\mathbb{R})$, and the trace operation $\mathfrak{tr}$ defines the bilinear form used to contract generators. To describe the dynamics of dilaton gravity in this setting, we utilize the basis of $sl(2,\mathbb{R})$ generators ${\tt L}_{i}$ with indices $i=0,\pm 1$, which obey the following commutation relations:
$\left[{\tt L}_{i},{\tt L}_{j}\right]$ $=$ $(i-j){\tt L}_{i+j}.$ (12)
To proceed with explicit computations, we adopt a matrix representation for the generators:
${\tt L}_{-1}=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}\quad,\quad{\tt L}_{0}=\frac{1}{2}\begin{pmatrix}-1&0\\ 0&1\end{pmatrix}\quad,\quad{\tt L}_{1}=\begin{pmatrix}0&-1\\ 0&0\end{pmatrix}\;$. (13)
With this representation, the only non–vanishing components of the invariant bilinear form are given by $\mathfrak{tr}({\tt L}_{\mp 1}{\tt L}_{\pm 1})=-2$ and $\mathfrak{tr}({\tt L}_{0}{\tt L}_{0})=-\tfrac{1}{2}$. The action remains invariant under the gauge transformations:
$\delta_{\lambda}{\cal A}=d\lambda+\left[{\cal A},\lambda\right],\qquad\qquad\delta_{\lambda}{\cal X}=[{\cal X},\lambda],$ (14)
Here, $\lambda$ is a gauge parameter taking values in the $sl(2,\mathbb{R})$ Lie algebra. These gauge transformations define the residual symmetries of the theory that preserve the asymptotic structure of the gauge fields, and they also determine how boundary symmetries emerge from bulk gauge invariance within the BF formulation. These elements provide the foundation for identifying the asymptotic symmetry algebra, which in the case of $sl(2,\mathbb{R})$ corresponds to the Virasoro algebra. The presence of the dilaton field plays a pivotal role in determining which symmetries are preserved and which are broken at the boundary. The equations of motion are given by
$\mathcal{F}=d{\cal A}+{\cal A}\wedge{\cal A}=0,\qquad\qquad d{\cal X}+\left[{\cal A},{\cal X}\right]=0.$ (15)
More importantly, these gauge transformations determine the residual symmetries that persist after imposing the radial gauge near the asymptotic boundary of $\mathrm{AdS}_{2}$. Consequently, they provide the foundational structure for identifying the asymptotic symmetry algebra. The inclusion of the dilaton field $X$ is especially significant in this context: it couples to the gauge connection at the boundary and introduces dynamical constraints that restrict the admissible gauge transformations. This interaction results in a partial breaking of the full infinite–dimensional affine symmetry algebra, such as $\mathfrak{sl}(2,\mathbb{R})_{k}$ , to its corresponding finite–dimensional subalgebra, namely $\mathfrak{sl}(2,\mathbb{R})$. In this light, equations (14) and (15) do more than define local gauge redundancy; they also encapsulate the physical mechanism through which symmetry breaking is dynamically induced by the dilaton field.
Adopting the radial gauge, the gauge connections in an asymptotically $\mathrm{AdS}_{2}$ spacetime can be expressed in the following factorized form:
${\cal A}$ $=$ $b^{-1}a\left(t\right)b+b^{-1}\mathrm{d}b,\quad{\cal X}=b^{-1}x\left(t\right)b,$ (16)
Here, the group element $b(\rho)$, which is state–independent and depends only on the radial coordinate, is chosen to take the form:
$b(\rho)=e^{\rho{\tt L}_{0}},$ (17)
This representation allows for a more general class of metrics that incorporate all $sl(2,\mathbb{R})$ charges. Importantly, as long as $\delta b=0$, the specific choice of $b$ does not affect the structure of the asymptotic symmetries. This freedom enables a more versatile formulation of the metric and requires boundary conditions that preserve such generality in the gravitational context. Additionally, within the radial gauge, the reduced connection $a(t)$ is a one-form valued in the $sl(2,\mathbb{R})$ Lie algebra and is independent of the radial coordinate, i.e., $a(t)=a_{t}(t)\,\mathrm{d}t$.
Our analysis focuses on the affine boundary conditions relevant to $sl(2,\mathbb{R})$ dilaton gravity. In doing so, we make use of the methodology developed in Grumiller:2016pqb ; Ozer:2017dwk ; Ozer:2019nkv ; Ozer:2021wtx ; Ozer:2024ovo to systematically derive the associated asymptotic symmetry algebra. Based on these techniques, the most general solution consistent with asymptotically $AdS_{2}$ geometries can be expressed in the following general metric form:
$\operatorname{d}\!s^{2}=\operatorname{d}\!\rho^{2}+2\,\mathcal{L}^{0}\operatorname{d}\!\rho\operatorname{d}\!\varphi+\Big(\left(e^{\rho}\mathcal{L}^{+}-e^{-\rho}\mathcal{L}^{-}\right)^{2}+\left(\mathcal{L}^{0}\right)^{2}\Big)\,\operatorname{d}\!\varphi^{2}\penalty 10000\ ,$ (18)
which bears resemblance to its well–known $AdS_{3}$ counterpart Grumiller:2016pqb ; Ozer:2017dwk ; Ozer:2019nkv ; Ozer:2021wtx ; Ozer:2024ovo ; Sammani:2025wtx . The dilaton field in this setting takes the general form:
$X=e^{\rho}\,\mathcal{X}^{+}+e^{-\rho}\,\mathcal{X}^{-}\penalty 10000\ .$ (19)
Accordingly, it becomes essential to specify affine $sl(2,\mathbb{R})$ boundary conditions that preserve this generalized metric structure.
3.3 Affine boundary conditions
The objective of this section is to formulate $sl(2,\mathbb{R})$ higher–spin $AdS_{2}$ dilaton gravity under affine boundary conditions. Our analysis proceeds by computing the asymptotic symmetry algebra associated with the most general (i.e., least restrictive) set of boundary data. We begin by introducing the $a_{t}$ component of the gauge connection, which takes values in the $sl(2,\mathbb{R})$ Lie algebra:
$a_{t}$ $=$ $\alpha_{i}\mathcal{L}^{i}{\tt L}_{i},$ (20)
where the coefficients satisfy $\alpha_{0}=-2\alpha_{\pm 1}=\frac{4}{k}$. The functions $\mathcal{L}^{i}$ are state-dependent and interpreted as dynamical boundary charges. Similarly, the dilaton field $x$ is expanded as
$x$ $=$ ${\cal X}^{i}{\tt L}_{i},$ (21)
which defines an additional set of three state–dependent functions ${\cal X}^{i}$. Our aim is to derive the resulting asymptotic symmetry algebra under affine boundary conditions through a canonical analysis. For this purpose, we systematically investigate all gauge transformations of the form (14) that preserve the boundary data encoded in equations (20) and (49).
At this stage, it is convenient to parameterize the gauge transformation in terms of the $sl(2,\mathbb{R})$ Lie algebra basis:
$\lambda=b^{-1}\left[\epsilon^{i}{\tt L}_{i}\right]b.$ (22)
The gauge parameter $\lambda$ thus contains three bosonic functions $\epsilon^{i}$, which are arbitrary functions defined on the boundary. We now focus on those gauge parameters that satisfy the transformation law (14). The corresponding infinitesimal variations of the dynamical fields are given by:
$2\delta_{\lambda}\mathcal{L}^{0}$ $=-\frac{k}{4}\partial_{t}\epsilon^{0}-\mathcal{L}^{+1}\epsilon^{-1}+\mathcal{L}^{-1}\epsilon^{+1},$ (23)
$\delta_{\lambda}\mathcal{L}^{\pm 1}$ $=\frac{k}{2}\partial_{t}\epsilon^{\mp 1}+\mathcal{L}^{0}\epsilon^{\mp 1}\mp\mathcal{L}^{\mp 1}\epsilon^{0},$ (24)
$\delta_{\lambda}\mathcal{X}^{0}$ $=2\big(\mathcal{X}^{+1}\epsilon^{-1}-\mathcal{X}^{-1}\epsilon^{+1}\big),$ (25)
$\delta_{\lambda}\mathcal{X}^{\pm 1}$ $=\mp\mathcal{X}^{0}\epsilon^{\pm 1}\pm\mathcal{X}^{\pm 1}\epsilon^{0}.$ (26)
As a final step in this construction, we define the canonical boundary charge $\mathcal{Q}_{a}[\lambda]$, which generates the gauge transformations of the connection field described in Eqs. (23)–(24). In parallel, we introduce the canonical boundary charge $\mathcal{Q}_{x}[\lambda]$, associated with the transformations of the dilaton field as specified in Eqs. (25)–(26).
The infinitesimal variation of these charges, which encodes the structure of the asymptotic symmetry algebra, is given by Banados:1994tn :
$\delta_{\lambda}\mathcal{Q}_{a}=\frac{k}{2\pi}\int\mathrm{d}t\;\mathfrak{tr}\left(\lambda\delta a_{t}\right),\quad\delta_{\lambda}\mathcal{Q}_{x}=\frac{k}{2\pi}\int\mathrm{d}t\;\mathfrak{tr}\left(\lambda\delta x\right).$ (27)
By performing a functional integration of the variations, we obtain the explicit form of the canonical boundary charges:
$\mathcal{Q}_{a}[\lambda]=\int\mathrm{d}t\;\left(\mathcal{L}^{i}\epsilon^{-i}\right),\quad\mathcal{Q}_{x}[\lambda]=\int\mathrm{d}t\;\left(\mathcal{X}^{j}\epsilon^{-j}\right).$ (28)
Having established both the infinitesimal gauge transformations and the associated canonical boundary charges, we now derive the asymptotic symmetry algebra using the standard procedure outlined in Blagojevic:2002du . The algebraic structure is encoded in the following relation:
$\delta_{\lambda}\digamma=\{\digamma,\mathcal{Q}_{a,x}[\lambda]\}$ (29)
[…]
$\mathcal{L}^{i}(\tau_{1})\mathcal{L}^{j}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\eta^{ij}}{\tau_{12}^{2}}+\frac{(i-j)}{\tau_{12}}\mathcal{L}^{i+j}$ (30)
$\mathcal{L}^{i}(\tau_{1})\mathcal{X}^{j}(\tau_{2})$ $\sim$ $\frac{(2i+j)}{\tau_{12}}\mathcal{X}^{i+j},$ (31)
where $\tau_{12}=\tau_{1}-\tau_{2}$. The coefficients $\eta^{ij}=\mathfrak{tr}({\tt L}_{i}{\tt L}_{j})$ define the components of the invariant bilinear form in the fundamental representation of the $sl(2,\mathbb{R})$ Lie algebra. An alternative and more compact representation of the operator product algebra can be written using index-free notation as follows:
$\mathfrak{\mathcal{\mathfrak{J}}}^{A}(\tau_{1})\mathfrak{\mathcal{\mathfrak{J}}}^{B}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\eta^{AB}}{\tau_{12}^{2}}+\frac{\mathfrak{\mathcal{\mathfrak{f}}}^{AB}_{\penalty 10000\ \penalty 10000\ \penalty 10000\ C}\mathfrak{\mathcal{\mathfrak{J}}}^{C}}{\tau_{12}}.$ (33)
Here, $\eta^{AB}$ denotes the trace matrix, and $\mathfrak{f}^{AB}_{\penalty 10000\ \penalty 10000\ C}$ are the structure constants of the underlying algebra, with indices running over $(A,B=0,\pm 1)$. In particular, one has $\mathfrak{f}^{ij}_{\penalty 10000\ \penalty 10000\ i+j}=(i-j)$. Accordingly, under the most relaxed set of boundary conditions for $sl(2,\mathbb{R})$ dilaton gravity, the resulting asymptotic symmetry algebra is identified with a single copy of the affine $\mathfrak{sl}(2,\mathbb{R})_{k}$ algebra.
The time–dependent profile of the dilaton field at the affine boundary introduces structural modifications that go beyond the classical asymptotic symmetry algebra of $sl(2,\mathbb{R})$ JT dilaton gravity. In the standard BF formulation, the asymptotic symmetry algebra is determined solely by gauge transformations of the $sl(2,\mathbb{R})$–valued connection $a_{t}$. However, once the $sl(2,\mathbb{R})$–valued dilaton field $x$ is treated as a dynamical degree of freedom, the set of residual symmetries becomes further constrained, giving rise to an enriched physical structure. In this extended framework, the symmetry algebra must incorporate not only the reparametrization modes but also new dynamical currents induced by the dilaton.
As a consequence, the full symmetry algebra necessitates the inclusion of abelian currents alongside the classical $\mathfrak{sl}(2,\mathbb{R})_{k}$ affine modes. This reveals that the affine boundary dynamics of JT gravity cannot be adequately described by the Schwarzian action alone, but instead call for a more general, multi-mode effective action. In this refined picture, the dilaton field plays a dual role: it determines both the geometric background and the extended algebraic structure realized at the boundary.
3.4 Conformal boundary conditions
In this section, our goal is to analyze the asymptotic symmetry algebra associated with Brown–Henneaux-type boundary conditions. This gauge choice is specifically designed to make the asymptotic symmetry structure manifest and to streamline the analysis of boundary dynamics. By decoupling the radial dependence, it effectively isolates the relevant asymptotic degrees of freedom, thereby facilitating the computation of conserved quantities and the identification of corresponding boundary theories.
This gauge plays a central role in maintaining the full asymptotic symmetry group and in determining the effective boundary action, particularly within the context of holographic duality. To proceed, we impose the Drinfeld–Sokolov highest weight gauge condition on the $sl(2,\mathbb{R})$-valued connection defined in Eq. (20), which further restricts the allowed functional form of the coefficients. Under this reduction, the fields are constrained as follows:
$\mathcal{L}^{0}=0,\quad\mathcal{L}^{-1}=\mathcal{L},\quad\alpha_{+1}\mathcal{L}^{+1}=1.$ (34)
where $\alpha_{-1}=\alpha$. It is important to note that these conformal boundary conditions correspond to the well–established Brown–Henneaux boundary conditions originally formulated in Brown:1986nw for $AdS_{2}$ gravity. Inspired by analogous developments in the context of three–dimensional gravity Perez:2016vqo ; Cardenas:2021vwo , we propose that the gauge connection and the dilaton field adopt the following specific structure:
$a_{t}$ $={\tt L}_{+1}+\alpha\mathcal{L}{\tt L}_{-1},$ (35)
$x$ $={\cal X}{\tt L}_{1}-{\cal X}^{\prime}{\tt L}_{0}+\bigg(\alpha\mathcal{L}{\cal X}+\frac{{\cal X}^{\prime\prime}}{2}\bigg){\tt L}_{-1}.$ (36)
Here, $\alpha$ acts as a scaling parameter whose exact value will be fixed at a later stage. We define two key functionals: the boundary charge $\mathcal{L}$ and the dilaton field $\cal X$}. With these constraints in place, we are now prepared to derive the conformal asymptotic symmetry algebra. As dictated by the Drinfeld–Sokolov reduction, the residual gauge parameter $\lambda$ depends on four independent functions, with $\epsilon\equiv\epsilon^{+1}$ serving as the primary functional degree of freedom. The gauge parameter then takes the following form:
$\lambda$ $=$ $b^{-1}\left[\epsilon{\tt L}_{1}-\epsilon^{\prime}{\tt L}_{0}+\bigg(\alpha\mathcal{L}\epsilon+\frac{\epsilon^{\prime\prime}}{2}\bigg){\tt L}_{-1}\right]b.$ (37)
The dilaton field $x$ is defined in the same form as the gauge parameter $\lambda$ in order to maintain consistency with the equations of motion and compatibility with the asymptotic symmetry group. This structural alignment allows the dilaton to behave as a coadjoint element in the gauge theory and to act as a stabilizing agent for the residual symmetries. Moreover, this choice ensures the well-posedness of the variational principle and enables the consistent derivation of the conserved Casimir function $C$. Thus, defining $x$ in the same form as $\lambda$ is crucial for preserving the asymptotic symmetry group, ensuring the consistency of the theory, and facilitating the formulation of boundary dynamics in a holographic context.
By substituting the gauge parameter $\lambda$ into the field transformation laws (14), the resulting infinitesimal gauge transformations are found to be:
$\delta_{\lambda}\mathcal{L}$ $=$ $\frac{1}{2\alpha}\epsilon^{{}^{\prime\prime\prime}}+\epsilon\mathcal{L}^{\prime}+2\mathcal{L}\epsilon^{\prime},$ (38)
$\delta_{\lambda}{\cal X}$ $=$ $\epsilon{\cal X}^{\prime}-\epsilon^{\prime}{\cal X}.$ (39)
Under the imposed boundary conditions, the corresponding canonical charges $\mathcal{Q}_{a}[\lambda]$ and $\mathcal{Q}_{x}[\lambda]$ remain integrable, thereby allowing $\mathcal{L}$ and $\mathcal{X}$ to be interpreted as conserved quantities. These gauge transformations are instrumental in uncovering the structure of the asymptotic symmetry algebra Banados:1998pi .
By analyzing the asymptotic symmetries, one can integrate the variations of the canonical boundary charges, $\delta_{\lambda}\mathcal{Q}_{a}$ as given in Eq. (38) and $\delta_{\lambda}\mathcal{Q}_{x}$ in Eq. (39), to arrive at the following expressions:
$\mathcal{Q}_{a}[\lambda]=\int\mathrm{d}\varphi\;\left(\mathcal{L}\epsilon\right),\quad\mathcal{Q}_{x}[\lambda]=\int\mathrm{d}\varphi\;\left(\mathcal{X}\epsilon\right).$ (40)
These canonical boundary charges naturally induce an operator product algebra at the asymptotic boundary. In contrast to the standard two–dimensional CFT framework, a one-dimensional CFT description can be obtained by effectively suppressing one of the coordinates. This is achieved by setting $z_{2}=0$ and identifying $z_{1}$ with the boundary coordinate $\tau$, or more simply by taking $z=\tau$:
$\mathcal{L}(\tau_{1})\mathcal{L}(\tau_{2})$ $\sim$ $\frac{{3k}}{{2\tau_{12}^{4}}}+\frac{2\mathcal{L}}{{\tau_{12}^{2}}}+\frac{\mathcal{L}^{\prime}}{\tau_{12}},$ (41)
${\cal L}(\tau_{1}){\cal X}(\tau_{2})$ $\sim$ $-\frac{{\cal X}}{{\tau_{12}^{2}}}+\frac{{\cal X}^{\prime}}{\tau_{12}}.$ (42)
From these relations, it follows that $\cal X$} transforms as a vector under boundary diffeomorphisms, while $\mathcal{L}$ behaves as a spin-two density. Throughout this analysis, the explicit dependence on the gauge parameter $\lambda$ has been omitted, as the boundary conditions no longer require it. In effect, the residual symmetries associated with the connection $a_{t}$ and the dilaton $x$ are interpreted as diffeomorphisms on the boundary circle, generated by the vector field $\epsilon$. It is thus observed that the asymptotic symmetry algebra of $sl(2,\mathbb{R})$ dilaton gravity with conformal boundary conditions is given by a single copy of the Virasoro algebra with central charge $c=3k$.
The conformal boundary profile of the dilaton field induces a structural extension analogous to the affine case, leading to an enriched version of the classical asymptotic symmetry algebra in $sl(2,\mathbb{R})$ JT gravity. In the conventional BF formulation, the asymptotic symmetry algebra is generated solely by gauge transformations of the $sl(2,\mathbb{R})$-valued connection $a_{t}$. However, once the dilaton field $x$ is promoted to a dynamical quantity, the residual symmetry subalgebra becomes more constrained, resulting in a more intricate physical structure. This generalized framework requires that the full symmetry algebra incorporate both the standard reparametrization modes and additional currents sourced directly by the dynamical behavior of the dilaton.
As a consequence, the boundary symmetry structure must necessarily include abelian extensions beyond the classical Virasoro algebra. This observation suggests that the Schwarzian action alone is insufficient to capture the complete boundary dynamics in the presence of a nontrivial dilaton configuration. Instead, a multi–component extended action is required to account for all dynamical features. In this context, the dilaton field assumes a dual role, encoding not only the background geometry but also the full algebraic structure realized at the boundary.
4 The $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ higher–spin dilaton supergravity
[…]
$S[{\cal X},{\cal A}]=\frac{k}{4\pi}\int_{{\cal M}}\mathfrak{str}[{\cal X}{\cal F}]+S_{\rm bdy},$ (43)
where $\mathcal{A}$ is a one–form superconnection with corresponding field strength ${\cal F}=\mathrm{d}{\cal A}+{\cal A}\wedge{\cal A}$, and $k$ denotes a coupling constant. The boundary term $S_{\rm bdy}$ is included to ensure a well-defined variational principle by enforcing suitable boundary conditions at $\partial\mathcal{M}$, which corresponds to the asymptotic boundary at infinite radius.
The dilaton field $\cal X$} is an algebra-valued scalar taking values in the Lie superalgebra $\mathfrak{osp}(1|2)$, just like the superconnection $\cal A$}. The supertrace operation $\mathfrak{str}$ provides the invariant bilinear form used to contract the generators of the superalgebra.
To describe the structure of the dilaton supergravity system, we now introduce the underlying $\mathfrak{osp}(1|2)$ superalgebra. It consists of bosonic generators $L_{n}$ with $n=0,\pm 1$ and fermionic generators $G_{p}$ with $p=\pm\tfrac{1}{2}$. These generators satisfy the following graded (anti-)commutation relations:
$\begin{split}\left[L_{m},L_{n}\right]&\,=\,(m-n)L_{m+n}\ ,\quad\left[L_{m},G_{p}\right]=\left(\frac{1}{2}m-p\right)G_{m+p}\ ,\\ \left\{G_{p},G_{q}\right\}&\,=\,-2L_{p+q}\ .\end{split}$ (44)
An explicit matrix representation of the $\mathfrak{osp}(1|2)$ generators can be written as:
$\begin{split}L_{1}\,=\,&\left[\begin{array}[]{c c|c}0&0&0\\ 1&0&0\\ \hline\cr 0&0&0\\ \end{array}\right]\ ,\quad L_{0}\,=\,\left[\begin{array}[]{cc|c}\frac{1}{2}&0&0\\ 0&-\frac{1}{2}&0\\ \hline\cr 0&0&0\\ \end{array}\right]\ ,\quad L_{-1}\,=\,\left[\begin{array}[]{c c|c}0&-1&0\\ 0&0&0\\ \hline\cr 0&0&0\\ \end{array}\right]\ ,\\ G_{\frac{1}{2}}\,=\,&\left[\begin{array}[]{c c|c}0&0&0\\ 0&0&1\\ \hline\cr-1&0&0\\ \end{array}\right]\ ,\quad G_{-\frac{1}{2}}\,=\,\left[\begin{array}[]{c c|c}0&0&1\\ 0&0&0\\ \hline\cr 0&1&0\\ \end{array}\right]\ .\end{split}$ (45)
In this basis, the invariant bilinear form is defined via the supertrace, and is given by:
$\mathfrak{str}\big(L_{n},L_{m}\big)=\begin{pmatrix}0&0&-1\\ 0&\frac{1}{2}&0\\ -1&0&0\end{pmatrix},\quad\mathfrak{str}\left\{G_{p}G_{q}\right\}=\begin{pmatrix}0&2\\ -2&0\end{pmatrix}$. (46)
where the supertrace $\mathfrak{str}$ of a matrix $\tt M$} is defined as
$\mathfrak{str}{\tt M}\equiv{\tt M}_{11}+{\tt M}_{22}-{\tt M}_{33}.$ (47)
4.1 $\mathcal{N}=1$ affine boundary conditions
The goal of this section is to construct an extended higher–spin $AdS_{2}$ framework by formulating it as an $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ dilaton supergravity theory defined with affine boundary conditions. Our analysis aims to determine the resulting asymptotic symmetry algebra under the most general boundary conditions consistent with this setup.
Following the discussion in the previous section, we adopt the principal embedding of $sl(2,\mathbb{R})$ into $\mathfrak{osp}(1|2)$ as a subalgebra. This embedding provides a natural foundation for implementing affine boundary conditions in asymptotically $AdS_{2}$ spacetimes. To proceed with this construction, we introduce the gauge connection in the following form:
$a_{t}=\alpha_{i}\mathcal{L}^{i}{\tt L}_{i}+\beta_{i}\mathcal{G}^{i}{\tt G}_{i}$ (48)
where the scaling relations among the parameters are given by $\alpha_{0}=-2\alpha_{\pm 1}=\mp 4\beta_{\pm\frac{1}{2}}$. As a consequence, the gauge connection contains five independent dynamical functions, denoted by $\mathcal{L}^{i}$ and $\mathcal{G}^{i}$, which are interpreted as the corresponding boundary charges.
The dilaton field $x$ is expanded in the same $\mathfrak{osp}(1|2)$ basis and takes the form:
$x$ $=$ $\mathcal{X}^{i}{\tt L}_{i}+\mathcal{Y}^{i}{\tt G}_{i}$ (49)
In this framework, we obtain eight additional state-dependent functions, ${\cal X}^{i}$ and ${\cal Y}^{i}$, arising from the dilaton field expansion. Our goal is to derive the asymptotic symmetry algebra associated with these affine boundary conditions via a canonical analysis. To this end, we systematically investigate all gauge transformations that preserve the structure of the boundary conditions defined above.
At this stage, it is convenient to express the gauge parameter in terms of the $\mathfrak{osp}(1|2)$ Lie superalgebra basis:
$\lambda=b^{-1}\left[\epsilon^{i}{\tt L}_{i}+\eta^{i}{\tt G}_{i}\right]b.$ (50)
Here, the gauge parameter consists of five bosonic functions, $\epsilon^{i}$ and $\eta^{i}$, which are arbitrary functions defined on the boundary. We now proceed to analyze those gauge parameters that satisfy the transformation law (14), and determine the corresponding infinitesimal gauge transformations for both the connection and the dilaton fields.
$\delta_{\lambda}\mathcal{L}^{\mp 1}$ $=-\frac{k}{2}\partial_{t}\epsilon^{\mp 1}\pm\eta^{\mp\frac{1}{2}}\mathcal{G}^{\mp\frac{1}{2}}\mp 2\mathcal{L}^{0}\epsilon^{\mp 1}\mp\mathcal{L}^{\mp 1}\epsilon^{0},$ (51)
$\delta_{\lambda}\mathcal{L}^{0}$ $=\frac{k}{4}\partial_{t}\epsilon^{0}-\frac{1}{2}\eta^{\frac{1}{2}}\mathcal{G}^{-\frac{1}{2}}+\frac{1}{2}\eta^{-\frac{1}{2}}\mathcal{G}^{\frac{1}{2}}-\mathcal{L}^{1}\epsilon^{-1}+\mathcal{L}^{-1}\epsilon^{1},$ (52)
$\delta_{\lambda}\mathcal{G}^{\mp\frac{1}{2}}$ $=\pm k\partial_{t}\eta^{\mp\frac{1}{2}}+2\eta^{\pm\frac{1}{2}}\mathcal{L}^{\mp 1}+2\eta^{\mp\frac{1}{2}}\mathcal{L}^{0}-\mathcal{G}^{\pm\frac{1}{2}}\epsilon^{\mp 1}\mp\frac{1}{2}\mathcal{G}^{\mp\frac{1}{2}}\epsilon^{0},$ (53)
$\delta_{\lambda}\mathcal{X}^{\mp 1}$ $=-2\eta^{\mp\frac{1}{2}}\mathcal{Y}^{\mp\frac{1}{2}}\mp\mathcal{X}^{\mp 1}\epsilon^{0}\pm\mathcal{X}^{0}\epsilon^{\mp 1},$ (54)
$\delta_{\lambda}\mathcal{X}^{0}$ $=-2\eta^{\frac{1}{2}}\mathcal{Y}^{-\frac{1}{2}}-2\eta^{-\frac{1}{2}}\mathcal{Y}^{\frac{1}{2}}+2\mathcal{X}^{1}\epsilon^{-1}-2\mathcal{X}^{-1}\epsilon^{1},$ (55)
$\delta_{\lambda}\mathcal{Y}^{\mp\frac{1}{2}}$ $=\pm\frac{1}{2}\eta^{\mp\frac{1}{2}}\mathcal{X}^{0}\mp\eta^{\pm\frac{1}{2}}\mathcal{X}^{\mp 1}\pm\mathcal{Y}^{\pm\frac{1}{2}}\epsilon^{\mp 1}\mp\frac{1}{2}\mathcal{Y}^{\mp\frac{1}{2}}\epsilon^{0}.$ (56)
As a final step in this construction, we define the canonical boundary charge $\mathcal{Q}_{a,x}[\lambda]$, which generates the gauge transformations presented in Eqs. (51)–(56). The variation of this charge, which encodes the structure of the asymptotic symmetry algebra, is given by Banados:1994tn :
$\delta_{\lambda}\mathcal{Q}_{a}=\frac{k}{2\pi}\int\mathrm{d}t\;\mathfrak{tr}\left(\lambda\delta a_{t}\right),\quad\delta_{\lambda}\mathcal{Q}_{x}=\frac{k}{2\pi}\int\mathrm{d}t\;\mathfrak{tr}\left(\lambda\delta x\right).$ (57)
Integrating this variation functionally yields the explicit expression for the canonical boundary charges:
$\mathcal{Q}_{a}[\lambda]=\int\mathrm{d}t\;\left(\mathcal{L}^{i}\epsilon^{-i}+\mathcal{G}^{i}\eta^{-i}\right),\quad\mathcal{Q}_{x}[\lambda]=\int\mathrm{d}t\;\left(\mathcal{X}^{i}\epsilon^{-i}+\mathcal{Y}^{i}\eta^{-i}\right).$ (58)
Having established both the infinitesimal gauge transformations and the canonical boundary charge, we are now in a position to derive the asymptotic symmetry algebra using the standard procedure outlined in Blagojevic:2002du . This is accomplished through the fundamental relation:
$\delta_{\lambda}\digamma=\{\digamma,\mathcal{Q}_{a,x}[\lambda]\}$ (59)
[…]
$\mathcal{L}^{i}(\tau_{1})\mathcal{L}^{j}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\eta^{ij}}{\tau_{12}^{2}}+\frac{(i-j)}{\tau_{12}}\mathcal{L}^{i+j},$ (60)
$\mathcal{L}^{i}(\tau_{1})\mathcal{G}^{j}(\tau_{2})$ $\sim$ $\frac{(\frac{i}{2}-j)}{\tau_{12}}\mathcal{G}^{i+j},$ (61)
$\mathcal{G}^{i}(\tau_{1})\mathcal{G}^{j}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\theta^{ij}}{\tau_{12}^{2}}+\frac{2}{\tau_{12}}\mathcal{L}^{i+j},$ (62)
$\mathcal{L}^{i}(\tau_{1})\mathcal{X}^{j}(\tau_{2})$ $\sim$ $\frac{(2i+j)}{\tau_{12}}\mathcal{X}^{i+j},$ (63)
$\mathcal{L}^{i}(\tau_{1})\mathcal{Y}^{j}(\tau_{2})$ $\sim$ $\frac{(\frac{i}{2}+j)}{\tau_{12}}\mathcal{Y}^{i+j},$ (64)
$\mathcal{G}^{i}(\tau_{1})\mathcal{X}^{j}(\tau_{2})$ $\sim$ $\frac{2}{\tau_{12}}\mathcal{Y}^{i+j},$ (65)
$\mathcal{G}^{i}(\tau_{1})\mathcal{Y}^{j}(\tau_{2})$ $\sim$ $\frac{(\frac{3i}{2}+\frac{j}{2})}{\tau_{12}}\mathcal{Y}^{i+j}.$ (66)
where $\tau_{12}=\tau_{1}-\tau_{2}$. In this formulation, $\eta^{ij}=\mathfrak{tr}({\tt L}_{i}{\tt L}_{j})$ and $\theta^{ij}=\mathfrak{tr}({\tt G}_{i}{\tt G}_{j})$ denote the invariant bilinear forms in the fundamental representation of the $\mathfrak{osp}(1|2)$ Lie superalgebra. The operator product algebra may be recast in a more compact, index-free form as follows:
$\mathfrak{\mathcal{\mathfrak{J}}}^{A}(\tau_{1})\mathfrak{\mathcal{\mathfrak{J}}}^{B}(\tau_{2})$ $\sim$ $\frac{\frac{k}{2}\eta^{AB}}{\tau_{12}^{2}}+\frac{\mathfrak{\mathcal{\mathfrak{f}}}^{AB}_{\penalty 10000\ \penalty 10000\ \penalty 10000\ C}\mathfrak{\mathcal{\mathfrak{J}}}^{C}}{\tau_{12}}.$ (67)
Here, $\eta^{AB}$ denotes the invariant trace matrix, while $\mathfrak{f}^{AB}_{\penalty 10000\ \penalty 10000\ C}$ represent the structure constants of the $\mathfrak{osp}(1|2)$ algebra, with indices $(A,B=0,\pm 1,\pm\tfrac{1}{2})$. These constants satisfy the relation $\mathfrak{f}^{ij}_{\penalty 10000\ \penalty 10000\ i+j}=(i-j)$. Under the most general (least restrictive) boundary conditions, the asymptotic symmetry algebra of $\mathfrak{osp}(1|2)$ dilaton supergravity is characterized by a single copy of the affine $\mathfrak{osp}(1|2)_{k}$ algebra.
In the context of $\mathfrak{osp}(1|2)$ JT dilaton gravity with affine boundary conditions, much like in the $sl(2,\mathbb{R})$ JT dilaton gravity under analogous assumptions, the asymptotic symmetry algebra is initially given by the infinite–dimensional $\mathfrak{osp}(1|2)_{k}$ algebra. However, the presence of a time-dependent dilaton field dynamically deforms this structure. As a result, even the expected preservation of the finite $\mathfrak{osp}(1|2)$ subalgebra within $\mathfrak{osp}(1|2)_{k}$ is partially broken. The dilaton acts as a stabilizing agent for the gauge superconnection and, through its boundary fluctuations, obstructs the full realization of the infinite–dimensional affine symmetry.
This symmetry breaking is attributed to the influence of the dilaton on free boundary excitations. Consequently, the residual symmetries encoded in the fields $a_{t}$ and $x$ correspond to boundary diffeomorphisms, generated respectively by the bosonic and fermionic components $\epsilon^{i}$ and $\eta^{i}$ of the gauge parameter.
4.2 $\mathcal{N}=1$ superconformal boundary conditions
In this section, we aim to analyze the asymptotic symmetry algebra associated with the Brown–Henneaux boundary conditions in the context of $\mathfrak{osp}(1|2)$ supergravity. To this end, we impose the Drinfeld–Sokolov highest weight gauge condition on the $\mathfrak{osp}(1|2)$-valued connection defined in Eq. (48), which further restricts the coefficient structure. This gauge fixing yields the following field constraints:
$\mathcal{L}^{0}=\mathcal{G}^{+\frac{1}{2}}=0,\quad\mathcal{L}^{-1}=\mathcal{L},\quad\mathcal{G}^{-\frac{1}{2}}=\mathcal{G},\quad\alpha_{+}\mathcal{L}^{+1}=1,$ (68)
and introduces the new scaling parameters $\alpha_{-1}=\alpha$ and $\beta_{-\frac{1}{2}}=\beta$.
It is worth noting that these conformal boundary conditions are in direct analogy with the well-known Brown–Henneaux boundary conditions initially developed in Brown:1986nw for $AdS_{3}$ gravity. Our approach draws on boundary conditions previously explored in the three–dimensional setting Perez:2016vqo ; Cardenas:2021vwo . In this spirit, we propose that the gauge connection and the dilaton field take the following form:
$a_{t}=$ ${\tt L}_{+1}+\alpha\mathcal{L}{\tt L}_{-1}+\beta\mathcal{G}{\tt G}_{-\frac{1}{2}},$ (69)
$x=$ $\mathcal{X}{\tt L}_{1}+\mathcal{Y}{\tt G}_{\frac{1}{2}}-\mathcal{X}^{\prime}{\tt L}_{0}+\left(\alpha\mathcal{G}\mathcal{X}-\mathcal{Y}^{\prime}\right){\tt G}_{-\frac{1}{2}}$ (70)
$+$ $\left(\alpha\mathcal{Y}\mathcal{G}+\beta\mathcal{L}\mathcal{X}+\frac{1}{2}\mathcal{X}^{\prime\prime}\right){\tt L}_{-1}$ (71)
After implementing the above procedures, we are now positioned to derive the conformal asymptotic symmetry algebra. As a consequence of the Drinfeld–Sokolov reduction, the residual gauge parameter $\lambda$ is effectively described by only four independent functions. These are denoted by $\epsilon\equiv\epsilon^{+1}$ and $\eta\equiv\eta^{+\frac{1}{2}}$, and are explicitly given by:
$\lambda=$ $b^{-1}\Bigg[\epsilon{\tt L}_{1}+\eta{\tt G}_{\frac{1}{2}}-\epsilon^{\prime}{\tt L}_{0}+\left(\alpha\mathcal{G}\epsilon-\eta^{\prime}\right){\tt G}_{-\frac{1}{2}}$ (72)
$+$ $\left(\alpha\eta\mathcal{G}+\beta\mathcal{L}\epsilon+\frac{1}{2}\epsilon^{\prime\prime}\right){\tt L}_{-1}\Bigg]b.$
By substituting this gauge parameter into the field transformation equation given in Eq. (14), we obtain the following infinitesimal gauge transformations:
$\delta_{\lambda}\mathcal{L}$ $=\frac{k\epsilon^{(3)}}{4}+2\mathcal{L}\epsilon^{\prime}+\epsilon\mathcal{L}^{\prime}-\frac{1}{2}\eta\mathcal{G}^{\prime}-\frac{3}{2}\mathcal{G}\eta^{\prime}$ (73)
$\delta_{\lambda}\mathcal{G}$ $=\frac{c\eta^{\prime\prime}}{3}+\epsilon\mathcal{G}^{\prime}+\frac{3\mathcal{G}\epsilon^{\prime}}{2}+2\eta\mathcal{L},$ (74)
$\delta_{\lambda}{\cal X}$ $=\epsilon\mathcal{X}^{\prime}-\mathcal{X}\epsilon^{\prime}-2\eta\mathcal{Y},$ (75)
$\delta_{\lambda}{\cal Y}$ $=\frac{\eta\mathcal{X}^{\prime}}{2}-\mathcal{X}\eta^{\prime}+\epsilon\mathcal{Y}^{\prime}-\frac{\mathcal{Y}\epsilon^{\prime}}{2}$ (76)
Under the specified boundary conditions, the canonical boundary charges $\mathcal{Q}_{a}[\lambda]$ are both well-defined and integrable, allowing the fields $\mathcal{L}$ and $\mathcal{G}$ to be interpreted as conserved charges. Likewise, the charges $\mathcal{Q}_{x}[\lambda]$ associated with the dilaton sector are also integrable, enabling a similar interpretation for $\mathcal{X}$ and $\mathcal{Y}$.
Importantly, these gauge transformations provide key insights into the structure of the asymptotic symmetry algebra Banados:1998pi . By analyzing the behavior of these asymptotic symmetries, one can determine the integral form of the variation of the canonical boundary charges,
$\mathcal{Q}_{a}[\lambda]=\int\mathrm{d}\varphi\;\left(\mathcal{L}\epsilon+\mathcal{G}\eta\right),\quad\mathcal{Q}_{x}[\lambda]=\int\mathrm{d}\varphi\;\left(\mathcal{X}\epsilon+\mathcal{Y}\eta\right).$ (77)
These canonical boundary charges serve as a fundamental tool in characterizing the asymptotic operator product expansion at the conformal boundary:
$\mathcal{L}(\tau_{1})\mathcal{L}(\tau_{2})\sim$ $\frac{{\frac{3k}{2}}}{{\tau_{12}^{4}}}+\frac{2\mathcal{L}}{{\tau_{12}^{2}}}+\frac{\mathcal{L}^{\prime}}{\tau_{12}},$ (78)
$\mathcal{L}(\tau_{1})\mathcal{G}(\tau_{2})\sim$ $\frac{\frac{3}{2}\mathcal{G}}{{\tau_{12}^{2}}}+\frac{\mathcal{G}^{\prime}}{\tau_{12}},$ (79)
$\mathcal{G}(\tau_{1})\mathcal{G}(\tau_{2})\sim$ $\frac{{2k}}{{\tau_{12}^{3}}}+\frac{2\mathcal{L}}{\tau_{12}}.$ (80)
[…]
With regard to the dilaton sector, the operator products involving $\mathcal{X}$ and $\mathcal{Y}$ reveal additional structural features:
${\cal L}(\tau_{1}){\cal X}(\tau_{2})\sim$ $-\frac{{\cal X}}{{\tau_{12}^{2}}}+\frac{{\cal X}^{\prime}}{\tau_{12}},$ (81)
${\cal L}(\tau_{1}){\cal Y}(\tau_{2})\sim$ $-\frac{\frac{1}{2}{\cal Y}}{{\tau_{12}^{2}}}+\frac{{\cal Y}^{\prime}}{\tau_{12}},$ (82)
$\mathcal{G}(\tau_{1}){\cal X}(\tau_{2})\sim$ $-\frac{2{\cal Y}}{{\tau_{12}^{2}}},$ (83)
${\cal G}(\tau_{1}){\cal Y}(\tau_{2})\sim$ $-\frac{{\cal X}}{{\tau_{12}^{2}}}+\frac{{\cal X}^{\prime}}{2\tau_{12}}.$ (84)
These structures provide a natural framework for organizing the asymptotic symmetries in $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ dilaton supergravity and play a central role in determining the boundary dynamics in the presence of symmetry breaking. While the relations described above correspond to the undeformed $\mathcal{N}=1$ superconformal algebra, the inclusion of dilaton fields $\cal X$} and $\cal Y$} in the extended $\mathfrak{osp}(1|2)$ model modifies this structure. These fields contribute additional terms to the algebra through boundary constraints, resulting in symmetry breaking and a deformation of the standard commutation relations. The deformed structure retains the essential features of the $\mathcal{N}=1$ superconformal algebra, while also encoding the dynamical influence of the dilaton sector.
[…]
$\cal G$} $\frac{3}{2}$ $\cal Y$} $\frac{1}{2}$
Table 1: superconformal spins for superconformal charges and related dilatons.
In the context of $sl(2,\mathbb{R})$ JT dilaton gravity, as well as in its $\mathfrak{osp}(1|2)$ extension under conformal boundary conditions, the asymptotic symmetry algebra at the boundary is initially identified with the infinite–dimensional $\mathcal{N}=1$ superconformal algebra. However, the presence of a time–dependent dilaton field alters the algebraic structure of the asymptotic symmetry. As a result, even the expected preservation of the finite $\tt{O}\tt{S}p(1|2)$ subgroup within the full $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ symmetry is dynamically reduced, and the remaining infinite–dimensional symmetry components are broken.
To illustrate the mechanism of symmetry breaking induced by the dilaton supermultiplet, we present a simplified toy model based on a static, partially truncated configuration. Consider the following profile for the $\mathfrak{osp}(1|2)$–valued dilaton field:
$\mathcal{X}^{0}=\text{const},\quad\mathcal{X}^{\pm 1}=0,\quad\mathcal{Y}^{\pm\frac{1}{2}}=0.$
This choice retains only the $\mathcal{X}^{0}$ component while suppressing all fermionic modes $\mathcal{Y}^{\pm\frac{1}{2}}$ and the bosonic charges $\mathcal{X}^{\pm 1}$. Under the gauge transformations generated by the full affine algebra (Eqs. (51)–(56)), one finds that such a configuration leads to a restricted set of residual symmetries. In particular, the variations $\delta_{\lambda}\mathcal{Y}^{\pm\frac{1}{2}}$ and $\delta_{\lambda}\mathcal{X}^{\pm 1}$ vanish identically, while $\delta_{\lambda}\mathcal{X}^{0}$ survives only for $\epsilon^{0}$ and a subset of $\eta$–parameters. As a consequence, the boundary charges associated with broken modes cease to transform non–trivially and are effectively projected out of the asymptotic symmetry algebra. This demonstrates, at a minimal level, how specific dilaton backgrounds can dynamically suppress certain symmetry generators, resulting in a reduced realization of the full $\mathfrak{osp}(1|2)_{k}$ structure. This kind of toy model construction, where a specific field configuration is used to analyze residual symmetry algebras, is also employed in the context of AdS2 boundary conditions Grumiller:2017qao .
The dilaton field plays the role of a stabilizer for the $\mathfrak{osp}(1|2)$ gauge connection, interacting with it through boundary fluctuations and thereby obstructing the complete realization of the full infinite–dimensional $\mathcal{N}=1$ $\mathfrak{osp}(1|2)$ symmetry. This symmetry reduction originates from the fluctuating behavior of the dilaton at the boundary. Table $1$ summarizes the superconformal charges and their corresponding dilaton spin values. Consequently, the residual symmetries encoded in the fields $a_{t}$ and $x$ manifest as diffeomorphisms on the boundary circle, generated by the vector fields $\epsilon$ and $\eta$.
We therefore identify the asymptotic symmetry algebra as the classical extended $\mathcal{N}=1$ superconformal algebra, rather than its quantum counterpart. This classical algebra arises naturally in our model and is consistent with the symmetry–breaking mechanism induced by the dilaton. Up to this point, our discussion has remained within the classical regime of the extended $\mathcal{N}=1$ superconformal symmetry. The quantum version of the theory, however, has already been presented in Eqs. (4)–(10). It is natural to ask whether the quantum theory can also be formulated in terms of a classical extended superconformal symmetry, or whether it necessitates a distinct quantum realization. While this remains a nontrivial question, preliminary evidence suggests that the quantum structure may be realized consistently, preserving the essential features of the classical algebra.
It is important to emphasize that the conformal boundary conditions considered in this section are not merely a technical simplification, but rather encode the infrared dynamics of the boundary theory. While the affine boundary data support a richer set of boundary modes—including commuting currents and higher–spin extensions—the conformal reduction effectively isolates the pseudo–Goldstone mode associated with the breaking of reparametrization symmetry. This reduction aligns with the low–energy behavior of dual SYK–like systems and captures the dominant physical degree of freedom via the Schwarzian effective action. Therefore, the transition from affine to conformal boundary conditions should be viewed as a flow from a general UV–complete description to an IR–dominated effective theory.
4.3 Scope and physical context
The present work is conceived as the $\mathcal{N}=1$ supersymmetric extension of our earlier analysis of $\mathfrak{sl}(3,\mathbb{R})$ JT–gravity Ozer:2025bpb . In that framework, asymptotic symmetry structures were studied from a bulk perspective, without recourse to Schwarzian–type boundary actions. The method focused on identifying residual gauge transformations and symmetry breaking patterns arising directly from the bulk gauge fields, providing a non–boundary formulation of the ASA.
In this supersymmetric extension, the same philosophy is maintained: we do not attempt to introduce a new effective boundary action, nor do we claim to formulate an alternative physical regime. Rather, we aim to generalize the bulk gauge–theoretic approach to the supersymmetric case, governed by the $\mathfrak{osp}(1|2)$ superalgebra.
The phrase “beyond the Schwarzian regime” is therefore not intended to suggest a new physical theory, but simply refers to a methodological shift away from boundary action–based analyses. This approach is rooted in the structure of gauge symmetries and their asymptotic realization, consistent with the strategy we established in the non–supersymmetric ${\tt S}{\tt L}(3,\mathbb{R})$ setting.
In this sense, while our construction remains algebraic in scope, the extended bulk symmetry breaking mechanism within the $\mathfrak{osp}(1|2)$ framework may serve as a foundation for future developments involving boundary dynamics, supersymmetric observables, or holographic models that transcend the traditional Schwarzian description. We believe that this perspective may complement ongoing efforts to understand JT supergravity beyond its standard boundary effective action paradigm.
From a physical standpoint, the dilaton-induced symmetry breaking mechanism analyzed in this work can be expected to influence observables in putative holographic dual theories. Specifically, the reduction of the full affine $\mathfrak{osp}(1|2)_{k}$ symmetry to its residual finite-dimensional subalgebra—together with the emergence of abelian commuting modes—modifies the spectrum of conserved charges at the boundary. This structure suggests a constrained class of boundary excitations, possibly leading to nontrivial selection rules or effective couplings in dual quantum mechanical models. While an explicit construction of the dual theory lies beyond the scope of this paper, we anticipate that the algebraic features derived here may provide structural constraints on supersymmetric SYK–like models or other one–dimensional superconformal systems. Further investigation of these physical implications remains an important direction for future work.
We conclude this section by briefly commenting on the quantum aspects alluded to above. While the central extension of the superalgebra emerges naturally from the classical coadjoint orbit analysis, potential operator ordering ambiguities can affect the precise realization of the symmetry generators at the quantum level. Addressing such ambiguities may require a more systematic treatment—such as path integral quantization or BRST analysis—which lies beyond the scope of this work. Nonetheless, our current results offer a robust classical foundation for such future quantum investigations, particularly in supersymmetric generalizations of the Schwarzian theory.
As a future direction, it would be interesting to explore quantum extensions of the symmetry breaking mechanisms presented here, especially in the context of super–Schwarzian dynamics or SYK–like boundary theories. The toy model introduced in Section 4.2 may also provide a useful testing ground for probing correlation functions and operator spectra in dual descriptions. We hope that our results provide a useful platform for further developments in two–dimensional supergravity and its holographic applications.
Quantum Extensions and Ordering Effects.
Beyond the classical analysis, the incorporation of quantum aspects introduces subtleties that merit closer attention. One such feature is the appearance of central extensions, which in our setup arise naturally from the coadjoint orbit construction. In particular, the Kirillov–Kostant symplectic structure on the orbit allows for a well-defined symplectic form, from which the classical central term in the $\mathcal{N}=1$ superconformal algebra can be read off Grumiller:2017qao ; Ikeda:1993fh . However, when promoting this classical structure to the quantum level, further issues emerge.
Most notably, operator ordering ambiguities may affect the precise realization of the symmetry generators and their algebra. These ambiguities stem from the non-commutative nature of quantum operators, especially in systems with fermionic generators and nontrivial constraints, such as in the $\mathfrak{osp}(1|2)$ case. While the classical Poisson brackets yield a consistent algebra, the quantum commutators require careful normal ordering or regularization prescriptions to preserve associativity and supersymmetry.
To properly resolve these quantum-level challenges, several quantization approaches can be considered:
  • Coadjoint orbit quantization: constructs the Hilbert space by quantizing the orbit equipped with a symplectic form Grumiller:2017qao ; Ikeda:1993fh . This technique is geometrically transparent but faces challenges in the presence of superalgebraic structure and boundary conditions.
  • BRST quantization: systematically incorporates gauge symmetries and constraints into the quantum framework Ikeda:1993fh ; Grumiller:2013swa . It is particularly suitable for supergravity models and BF-type gauge theories but often trades off geometric intuition for algebraic rigor.
  • Path integral quantization: provides a functional integral formulation of the theory, enabling the study of partition functions and quantum correlation functions Engelsoy:2016xyb ; Cardenas:2018krd . For $\mathfrak{osp}(1|2)$ supergravity, this approach captures the boundary dynamics and fluctuations of super-Schwarzian modes.
While a detailed quantum treatment lies beyond the present scope, the formal framework established in this work offers a solid foundation for such investigations. In particular, the toy model and the algebraic symmetry breaking discussed in Section 4.2 provide concrete starting points for analyzing quantum deformations and the structure of dual one-dimensional superconformal systems.
5 Conclusions and Discussion
Two–dimensional gravity models, most notably JT gravity, where the dilaton plays an active, dynamical role, have become central in the study of low–dimensional holography and quantum gravity. In its $sl(2,\mathbb{R})$ formulation, JT gravity defines the asymptotic conditions of $\mathrm{AdS}_{2}$ spacetime and allows spacetime diffeomorphisms to be enhanced into an infinite–dimensional Virasoro symmetry. However, the presence of a time–dependent dilaton field induces a nontrivial deformation in this structure, dynamically breaking even the expected preservation of the ${\tt S}{\tt L}(2,\mathbb{R})$ subgroup within the full Virasoro algebra. This breakdown is directly linked to dilaton fluctuations at the boundary, which govern the emergent dynamics of physical quantities near the asymptotic region. Thus, the dilaton evolves from a mere geometric label into a core dynamical variable that reshapes the boundary metric structure itself.
This foundational mechanism is further developed in the context of an $\mathfrak{osp}(1|2)$ supergravity extension explored in this work, where symmetry breaking and symmetry extension are jointly realized within a unified gauge–theoretic formulation.
As a natural supersymmetric extension of classical $sl(2,\mathbb{R})$ JT gravity, the $\mathfrak{osp}(1|2)$ JT supergravity framework introduces a structurally richer theory that incorporates both higher–order gauge connections and a more intricate configuration of boundary symmetries. This is not merely a formal enlargement of the algebra via fermionic generators; rather, it entails a nontrivial extension of the dynamical phase space and the gauge sector itself.
Notably, the $\mathfrak{osp}(1|2)$ gauge fields exhibit a novel mechanism of symmetry breaking, one that does not ensure the full preservation of even the $\tt{O}\tt{S}p(1|2)$ subgroup, resulting in substantial modifications to the boundary dynamics. This extended algebraic framework plays a key role in exploring generalized holographic dualities, wherein extended symmetries affect thermodynamic quantities such as energy and entropy. The resulting theory provides fertile ground for investigating the interplay between supersymmetry, higher–spin fields, and low–dimensional holography.
In contrast to earlier formulations of $\mathfrak{osp}(1|2)$ dilaton supergravity, most of which rely on coadjoint orbit techniques or metric–based descriptions, our BF–theoretic approach provides a systematic derivation of the full $\mathcal{N}=1$ superconformal algebra. More importantly, it unveils novel dynamical effects induced by the dilaton fields that have no analogue within the $\mathfrak{sl}(2,\mathbb{R})$ framework. These effects arise naturally from the gauge–theoretic treatment and highlight the deeper physical implications of supersymmetric extensions, beyond their algebraic structure alone.
The primary distinction between $sl(2,\mathbb{R})$ and $\mathfrak{osp}(1|2)$–based JT gravity theories lies in the role of the dilaton field and its impact on the extension of asymptotic symmetry algebras. While the $sl(2,\mathbb{R})$ structure is sufficient to encode the isometric features of $\mathrm{AdS}_{2}$ geometry, the $\mathfrak{osp}(1|2)$ framework offers a richer mathematical landscape to investigate more intricate holographic dualities and the dynamical behavior of boundary physical quantities. In particular, $\mathfrak{osp}(1|2)$ provides a natural starting point for systematically extending the theory to higher–rank super Lie algebras and Chern–Simons–based formulations.
The findings of this study demonstrate that the asymptotic symmetry algebra is governed not only by the structural features of the gauge fields but also by the dynamical configurations of the dilaton field near the boundary. The dilaton is not an external matter sector, it is an intrinsic and structural part of the theory, taking values in the adjoint representation of the gauge algebra. Under suitable boundary conditions, its time–dependent components become dynamical and contribute to the asymptotic symmetry algebra in the form of an abelian ideal composed of mutually commuting modes. This extension is not the result of adding an external $U(1)$ symmetry but instead arises from internal degrees of freedom that become dynamically relevant at the boundary. Consequently, the commuting currents induced by the dilaton should be viewed as a natural substructure within the extended asymptotic symmetry algebra, consistent with a semi–direct sum structure. The boundary behavior of the dilaton therefore plays a pivotal role in enriching the algebra while preserving its internal coherence, revealing the significance of internal gauge dynamics in shaping the asymptotic structure of two–dimensional gravitational systems.
In conclusion, the presence of the dilaton field profoundly affects not only the boundary dynamics in both $sl(2,\mathbb{R})$ and $\mathfrak{osp}(1|2)$ JT gravity frameworks, but also the mechanisms of extension and breaking within their respective asymptotic symmetry algebras. These extended models serve as powerful tools for uncovering new symmetry structures tied to boundary conditions and significantly enrich the theoretical landscape of holographic duality. In particular, they provide a robust foundation for probing deeper into holographic constructions such as the AdS/CFT correspondence and the SYK model. Future investigations into higher–rank Lie group–based generalizations are expected to further advance our theoretical understanding in this domain.
As discussed in the Introduction, several recent studies have examined the connection between generalized JT gravity and higher–spin extensions, particularly in the context of deformed SYK models and enhanced symmetry structures. For instance, in Momeni:2024ygv , a deformed SYK model incorporating spin–$3$ operators is proposed as a potential holographic dual of generalized dilaton gravity. Our construction based on $\mathfrak{osp}(1|2)$, equipped with an explicit $\mathcal{N}=1$ superconformal symmetry and a controlled symmetry breaking framework, offers a concrete realization of this duality. In parallel, the works Momeni:2020tyt ; Momeni:2020zkx ; Momeni:2021jhk explore complementary approaches grounded in coadjoint orbit techniques, the quantization of extended superconformal algebras, and generalized boundary dynamics. While these efforts share similar motivations, our approach is distinguished by its BF–theoretic formulation, which systematically derives the asymptotic symmetry algebra and elucidates the role of dilaton-induced symmetry breaking. This perspective advances the theoretical understanding of how higher–spin structures can be coherently integrated into two–dimensional holographic frameworks.
In our model, the interplay between the dilaton and higher–spin fields at the boundary induces a breakdown of the asymptotic symmetry structure. The initially present infinite–dimensional $\mathcal{N}=1$ superconformal algebra becomes dynamically reduced to its $\tt{O}\tt{S}p(1|2)$ subgroup under the influence of the dilaton. This mechanism closely resembles Schwarzian–type symmetry breaking. However, beyond this complete reduction, the boundary fluctuations of the dilaton field can also support instances of symmetry extension. As a result, our framework exhibits a novel coexistence of both asymptotic symmetry breaking and symmetry enhancement. For a comprehensive summary of boundary theories and their corresponding symmetry structures, see Appendix B.
In conclusion, JT gravity, formulated as a two–dimensional BF theory, offers complementary theoretical frameworks for probing the geometric and physical origins of asymptotic symmetries and Schwarzian dynamics. The gauge symmetries inherent to BF theory, together with the algebraic structure of the target space, form the foundation of asymptotic symmetry algebras at the boundary. JT gravity, in this context, serves as a specific realization of such structures. In both formulations, the Schwarzian action emerges as an effective reduced model capturing the residual symmetries inherited from bulk gauge transformations. However, the dynamics governed by the Schwarzian action do not encompass the entirety of the gauge freedom, but rather reflect the behavior of a physically significant subset. Therefore, BF theory and JT gravity jointly provide a rich and unified approach to elucidating the full structure of boundary symmetries and their holographic implications.
An open question remains as to whether the $\mathfrak{osp}(1|2)$ JT gravity framework developed in this work admits a well–defined quantum mechanical dual, possibly in the spirit of an extended SYK–like model involving spin-$\tfrac{3}{2}$ operators or deformed reparametrization modes. In our earlier $\mathfrak{sl}(3,\mathbb{R})$ analysis Ozer:2025bpb , we demonstrated how boundary dynamics can go beyond the classical Schwarzian regime. In a similar vein, the present study does not aim to reproduce the super–Schwarzian Cardenas:2018krd structure, but rather to explore boundary dynamics that transcend it. In this context, pursuing a consistent generalization based on extended $\mathfrak{osp}(\mathcal{N}|2)$ BF theories with $\mathcal{N}>1$ provides a promising direction for uncovering richer asymptotic structures and novel symmetry–breaking mechanisms in future work.
In addition, we have clarified how the affine–to–conformal transition may be interpreted as an infrared truncation, consistent with the dominance of pseudo–Goldstone boundary modes. We also outlined how coadjoint orbit, BRST, and path integral quantization methods may further illuminate the quantum structure of this framework. These observations suggest multiple pathways for extending the present work toward a more complete dual description.
While our construction does not include an explicit boundary action, the structure of the residual $\mathfrak{osp}(1|2)$ symmetries obtained from the bulk gauge formulation suggests a connection to extended super–Schwarzian–type quantum mechanical models. We emphasize that our focus remains on the intrinsic gauge–theoretic realization of boundary symmetries, and not on deriving a holographic dual theory. Nevertheless, the algebraic structure revealed in our analysis may serve as a guiding framework for future efforts aiming to identify or construct boundary effective actions consistent with this extended supersymmetric ASA. Whether such a theory generalizes or deviates from the standard Schwarzian sector remains an open direction that we hope to explore in future work.
This residual symmetry structure thus serves as a partial bulk–boundary dictionary: while we do not construct an explicit boundary theory, the algebraic content of the $\mathfrak{osp}(1|2)$ BF formulation indicates a possible correspondence with a generalized super-Schwarzian boundary dynamics.
6 Acknowledgments
Authors are supported by The Scientific and Technological Research Council of Türkiye (TÜBİTAK) through the ARDEB 1001 project with Grant number 123F255.
Appendix A Notation Table
Symbol Description
${\tt L}_{i},{\tt G}_{i}$ $\mathfrak{osp}(1|2)$ superalgebra basis
$\mathcal{L}^{i},\mathcal{G}^{i}$, $\mathcal{X}^{i},\mathcal{Y}^{i}$ affine boundary charges and dilaton components
$\epsilon^{i},\eta^{i}$ affine boundary variation parameters
$\mathcal{L},\mathcal{G}$, $\mathcal{X},\mathcal{Y}$ superconformal boundary charges and dilaton components
$\epsilon,\eta$ superconformal boundary variation parameters
Table 2: Notation Table
Appendix B Summary of Boundary Theories
Dilaton gravity theory Boundary fields Asymptotic symmetry algebra
$\mathfrak{sl}(2,\mathbb{R})$-BF affine $\mathcal{L}^{i},\mathcal{X}^{i}$ Extended affine $\mathfrak{sl}(2,\mathbb{R})_{k}$
$\mathfrak{sl}(2,\mathbb{R})$-BF conformal $\mathcal{L},\mathcal{X}$ Extended $\mathcal{W}_{2}$
$\mathfrak{osp}(1|2)$-BF affine $\mathcal{L}^{i},\mathcal{G}^{i},\mathcal{X}^{i},\mathcal{Y}^{i}$ Extended affine $\mathfrak{osp}(1|2)_{k}$
$\mathfrak{osp}(1|2)$-BF superconformal $\mathcal{L},\mathcal{G},\mathcal{X},\mathcal{Y}$ Extended $\mathcal{N}=1$ superconformal
Table 3: Boundary dynamics and asymptotic symmetries.